OpenStudy (anonymous):

16x – 3(4x + 7) = 6x –(2x + 21)what does x=?

OpenStudy (bahrom7893):

multiply out the parenthesis: 16x - 12x - 21 = 6x - 2x - 21 Add 21 to both sides: 16x - 12x = 6x - 2x Move all x's to one side: 16x - 12x - 2x - 6x = 0 x = 0.. are u sure the question is right?

OpenStudy (anonymous):

yes but it just doesnt make sense right!!!!!

OpenStudy (bahrom7893):

yeah.. but that is the right answer, x = 0

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

how

OpenStudy (anonymous):

I'm wrong but it end up being 1=1

OpenStudy (anonymous):

no oi just tried it and your right

OpenStudy (anonymous):

4x=4x so that means 4/4 =x/x which means 1=1

OpenStudy (anonymous):

or x=x

OpenStudy (anonymous):

Solve the system of equations using elimination. x-3y=-4 and 2x+6y=5

OpenStudy (anonymous):

Which means you can plug in any value for x and the equation holds true

OpenStudy (anonymous):

and how did u get it

OpenStudy (anonymous):

how did i get what?...

OpenStudy (bahrom7893):

for the first equation there are infinite number of solutions.. it's not just 0 or 1

OpenStudy (bahrom7893):

x-3y=-4 2x+6y=5 x = -4 + 3y 2(-4+3y)+6y = 5

OpenStudy (anonymous):

yeah thats why i said you can plug in any value for x and the equation holds true

OpenStudy (bahrom7893):

sorry didnt see that..

OpenStudy (anonymous):

its all good

OpenStudy (bahrom7893):

x = -4 + 3y 2(-4+3y)+6y = 5 x = -4 + 3y -8 + 6y + 6y = 5

OpenStudy (bahrom7893):

x = -4 + 3y 12y = 13 x = -4+3y y = 13/12

OpenStudy (bahrom7893):

x = -4+3y y = 13/12 x = -4 + 13/4 = -16/4 + 13/4 = -3/4 y = 13/12